OpenStudy (anonymous):

lim_(x->pi/2) (x-(pi/2))tanx ok so, i applied L'Hopital's Rule and got lim_(x->pi/2) tanx+(x-pi/2)sec^2(x) Now, plugin pi/2 and I got 1. The answer is -1 WolfRam said it was -1 too, but couldn't show me the steps. Thanks in advance :)

OpenStudy (anonymous):

lim as x --> pi/2 of { (x-pi/2) / (cosx/sinx) }

OpenStudy (anonymous):

that doesn't help...

OpenStudy (dumbcow):

When applying L'Hopitals Rule I believe you need to form a fraction rewrite expression \[(x-\pi/2)\tan x = \frac{x- \pi/2}{\cot x}\] \[\frac{d}{dx} \cot x = -\frac{1}{\sin^{2} x}\] \[\rightarrow \lim_{x \rightarrow \pi/2} \frac{1}{-\frac{1}{\sin^{2} x}} = -\sin^{2} x = -1\]

OpenStudy (anonymous):

how did you get \[(x-\pi/2)/\cot(x)\]?

OpenStudy (dumbcow):

from the fact that \[\frac{1}{\cot x} = \tan x\]

OpenStudy (dumbcow):

\[\cot x = \frac{1}{\tan x}\]

OpenStudy (anonymous):

ahh, ok. Thanks

OpenStudy (dumbcow):

your welcome